kekeman:

A flag maker is designing a new triangular flag. The angle of depression from the top of the flagpole is 30°. The height of the flag will be 15 inches and the bottom side will measure 26 inches. Part A) What must be the length of the top side of the flag? Round the answer to the nearest hundredth. Part B) What will be the area of the flag? Round the answer to the nearest hundredth.

Lyraleii:

The angle of depression from the top of the flagpole is 30° Using the tangent function, we have: tan(30°) = 15/x x = 15/tan(30°) To find the area of the flag, we can use the formula for the area of a triangle: Area = (1/2) * base * height the height of the flag is 15 inches and the length of the top side of the flag is x

KyledaGreat:

We can solve this problem using trigonometry. Let's call the length of the top side of the flag "x". Then, we can use the tangent function to find x. Part A: tan(30°) = 15/x x = 15/tan(30°) x ≈ 25.98 Therefore, the length of the top side of the flag must be approximately 25.98 inches. Part B: The area of a triangle can be found using the formula A = 1/2 * base * height. In this case, the base is 26 inches and the height is 15 inches, so we can substitute these values into the formula: A = 1/2 * 26 * 15 A = 195 Therefore, the area of the flag will be approximately 195 square inches.

kekeman:

Okay thank you both thank you!!!!!

surjithayer:

 \[\frac{ 15 }{x }=\sin 30=\frac{ 1 }{ 2}\] x=15 *2=30 in area\[=\frac{ 1 }{ 2 }\times 15 \times 26=195~in^2\]

kekeman:

@surjithayer wrote:
Created with Raphaël30602615x30Reply Using Drawing \[\frac{ 15 }{x }=\sin 30=\frac{ 1 }{ 2}\] x=15 *2=30 in area\[=\frac{ 1 }{ 2 }\times 15 \times 26=195~in^2\]
Thank you!!!